Prev Next

SW&IT Meeting Minutes 2018-10-15

Attendees: David, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

  • board meeting

Previous Action Items

  • xxx

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s)Sorted ascending Primary Secondary Week
ALH Igor(1/2)
Matthieu (2)
1
AzCam
AC (2)
Petr (2)
2
DIMM
AC (4)
Petr (2)
2
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
GCS
Petr (4)
Matthieu (2)
1
INDI Igor (1)
Matthieu (2)
3
LBC
Yang (3)
Igor(1/2) 2
MODS
Yang (2)
Igor (1) 2
OAC
AC (4)
Petr (4)
1
OVMS
Petr (4)
Matthieu (2)
3
particle monitors AC Petr 4
Status Server Stephen (2)
Matthieu (2)
2
Telem-VIZ
Yang (1)
Stephen (1) 3
Telemetry
Petr (4)
Matthieu (2)
2
Telemetry cleanup Stephen (1) Petr 3
Weatherstation
AC (2)
Petr (2)
2

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • FOS Wiki - ingest ARGOS/Arcetri
  • Magic laptop clone/vm - news from Microgate?
  • SOUL/FLAO Oct18 in lab then telescope
  • ARGOS/ci - add test, re-enable 'monit'
  • ARGOS deploy at telescope (Gustavo email)
  • Tucson CR, UPS/power plan
  • LBTI/shark+ilocater IIF/INDI
  • Vsphere upgrade, capture in plan
  • Luci web-io cold heads state in telemetry for vibration
  • Luci H/O doc comments from RR

This week in github

Team reports

YZ

  • Web front programming learning and prototyping:
    • Google Polymer 3 + Web components learning and prototyping.
    • Setting up a new Fedora 28 VM for dev use.
  • LBC/MODS documents reading/preparation.

AC

Last Week
  • ARGOS deployment
    • PR to add libPvAPI
    • updated user manual
    • successfully builds in Jenkins and Docker environments
  • Arcetri FOSWiki
    • Data tar file corrupt?

This Week
  • Vacation Monday
  • ARGOS deployment
    • configure build with ICS and TaN tests in Jenkins and Docker
    • deploy to OPS machines
  • Acetri FOSWiki
    • ask for new tarball

MB

  • management OC, WPC
  • jenkins-ci on going: epics/extensions/zero-c ice/python-xxx
  • review litHtml/Element, not quite there yet. Wait 1.0
  • HO, review INDI, sent questions to KS
  • skytech LBC update, contacted Fernando, no reply yet
  • review Guido PBP requirements document, talk to XY
  • discuss LBTI rack-space, network. rack-space=RR/JC. Network and IIF still unclear.
  • review Shark/ilocater interface
next:
  • review LBTI/new instruments interface - AO
  • new candidate phone interviews, post ad on select newsgroup
  • litHtml continued
  • sci-arch meeting prep

XZ

Last week
  • AGW connection test for SOUL
  • Read power backplane requirement document
  • Discuss the PBP document with Matthieu
  • Wednesday Thursday and Friday off the office
This week
  • Work on power backplane requirement document?
  • support the on-going AO work

SH

  • wiki to foswiki
  • AGW in lab Alessandro
  • LBTI problems with SOUL network
  • Find bandwidth hog of mountain network
  • Docker registry
This week
  • Docker test environment
  • ARGOS wiki import
  • cisrv01 RPM build
  • LBTI host/VLAN changes for Paul G.
  • Replace Xianyu's machine
  • Mark W. ssh issues to "vela".
  • Virtualize abell.as.arizona.edu (?) need to talk to John H.

PK

  • OVMS, GCS, Telemetry
  • finished reading TPK User Guide
  • IT 7387 - LBC NS fix
  • SX actuators problems
This
  • SX accuators troubleshooting
  • finish IT 7387
  • GCS

DE

Last week:
  • Broke archive-->GDrive backup into chunks to lower load, seems to be going steadily
  • Did a badblocks scan on the drive to make sure it was in good health. Results were good
  • AGw computer work
  • Reimaged with two methods, both of which still showed boot errors
    • Discovered that SATA cable was bad
  • Ordered replacement cables from reputable seller, should arrive today
  • 2x drive failure in Tucson Synology
    • Same volume, thankfully it was RAID-6
    • Drives have been replaced with spares and array is rebuilding
    • Failed drives sent for warranty replacement
This week:
  • Yang requested a new Fedora 28 VM for development
  • Button up AGw computer when the new SATA cable arrives
  • Draw up server room power maps
  • Work with Jim re: Autodesk software installs

KS

Last week
  • Updates to OVMS troubleshooting page to match the updated kernel (no longer 3 processes, 12 boards, LBTI is default instrument, etc.)
  • Last minute notes for LBC to YZ/IS as I was cleaning up.
  • Clean up the server of most of the uninteresting files.
This week
  • Pack up the remote server, monitors, phone, etc. and send to LBTO.

IS

Last week:
  • LUCI
    • Wrapping up RTD improvements.
    • Testing LUCI
  • OT
    • Meeting about AFC changes on the OT and how to tackle it.
  • Documentation
    • INDI/MODS only some reading
This week
  • Documentation:
    • INDI/ALH/LBC/MODS
  • OT
    • Meeting about AFC changes on the OT and how to tackle it
Topic revision: r17 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback